# Backpropagation-free Spiking Neural Networks with the Forward-Forward Algorithm

### v1

@@ 3 Proposed Method
MNIST F-MNIST CIFAR-10 N-MNIST SHD
Dataset (train/test) $60\text{k}/10\text{k}$ $60\text{k}/10\text{k}$ $50\text{k}/10\text{k}$ $60\text{k}/10\text{k}$ $8156/2264$
Input neurons $784$ $784$ 3072 $2312$ $700$
Dataset classes $10$ $10$ $10$ $10$ $20$
Epochs $300$ $300$ $300$ $300$ $500$
Learning rate $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
Batch size $4096$ $4096$ $4096$ $4096$ $4096$
Time steps $T$ $10$ $10$ $10$ $10$ $10$
$\alpha$ (in loss) $0.6$ $0.6$ $0.6$ $0.6$ $0.6$
Neurons’ threshold $1.0$ $1.0$ $1.2$ $1.0$ $5.0$
Neurons’ decay rate $0.99$ $0.99$ $0.8$ $0.9$ $0.9$
In the inference phase, due to the absence of an output layer, implementing conventional classification is not possible. Hence, a method should be adopted that uses layer information to classify test data. For each test sample, all possible labels in the dataset are applied to the sample once individually, and each of them passes to the trained neural network. After that, the sum of all goodness values of all layers is calculated, and the label of the sample that produces the highest sum of goodness values is considered the sample class.
@@
Table 2: Performance of our SNN model comparison to existing SNN models (* denotes self-implementation, denotes data augmentation, and β denotes learnable membrane time constants).
@@
F-MNIST Proposed model 784-1000-1000 LIFβ FF 89.82
@@
Taylor et al. [39] 784-1000-10 LIFβ BP 89.05
Neftci et al. [4] 784-1000-10 LIFβ BP 89.93
CIFAR-10 Proposed model 3072-2000-2000 LIF FF $51.05$
Neftci et al. [4]* 3072-2000-2000-10 LIF BP 47.82
@@
Neftci et al. [4] 700-300-20 LIFβ BP 70.81
Cramer et al. [36] 700-128-20 recurrent LIF BP 71.40
Table 3: Performance of our SNN model comparison to existing ANN implementation of the Forward-Forward algorithm
@@
CIFAR-10 Proposed model 3072-3072-3072 SNN 54.04
@@
4 Experiments
4.1 Experimental Setup
The experiments were conducted using a PyTorch-based implementation [30], leveraging the snntorch library [31] for spiking neuron models and surrogate gradients. To evaluate the proposed method, five datasets were utilized: MNIST [32], Fashion-MNIST (F-MNIST)[33], CIFAR-10 [34], Neuromorphic-MNIST (N-MNIST) [35], and Spiking Heidelberg Digits (SHD) [36]. The MNIST dataset comprises 70,000 grayscale images of hand-written digits, each sized 28×28 pixels. The dataset is divided into 60,000 samples for training and 10,000 for testing, covering digit classes from 0 to 9. Its variation, the F-MNIST dataset, retains the same number of samples and classes but replaces hand-written digits with images of fashion items such as T-shirts and shoes, providing more complex features for classification tasks. The CIFAR-10 dataset includes 60,000 color images, each sized 32×32 pixels with three color channels. It spans 10 object classes, with 50,000 samples allocated for training and 10,000 for testing. Unlike MNIST and Fashion-MNIST, which use grayscale images, CIFAR-10 focuses on more visually intricate static 2D color images. The N-MNIST dataset is derived from the MNIST dataset by applying a spiking conversion. This process involves recording digit samples displayed on an LCD monitor using an ATIS sensor mounted on a motorized pan-tilt unit, which introduces motion-based neuromorphic features while retaining the same classes and sample count as MNIST. Lastly, the SHD dataset is used, which includes high-quality studio recordings of spoken digits in both German and English. The original dataset contains 10,420 samples which are divided into 8,156 training and 2,264 testing samples spanning 20 classes.
4.2 Experimental Results and Analysis
Table 2 comprehensively compares spiking neural network models across five datasets: MNIST, FMNIST, CIFAR-10, N-MNIST, and SHD. The key columns detail the model name, architecture, neuron model type, and corresponding accuracy percentage, highlighting both our proposed model and various state-of-the-art models with fully connected architecture. In the MNIST dataset, our model achieves an impressive 98.69% accuracy using a LIFβ[37] neuron model and two hidden layers of 500 neurons architecture, outperforming other models. Compared to the FF-based models, Ororbia et al. use a deeper model with two hidden layers of 700 neurons, achieving 97.46% with LIF neurons, and Terres et al. achieve lower accuracy with 93.27%. Competing BP-based models from Neftci et al. [4], Comsa et al. [38], and Taylor et al. [39] use slightly different architectures and neuron models like IF or LIFβ, achieving comparable accuracies 97.87%, 97.90%, and 97.91%, respectively. However, our model outperforms these models, showcasing an improvement, potentially due to the specific architecture and neuron model used.
In the F-MNIST, our model achieves an accuracy of 89.82% with the same two hidden layers of 1000 neurons architecture and LIFβ neuron model. Compared to an FF-based model, Terres et al. use two hidden layers of 1400 neurons, achieving 85.68% that our model achieves higher accuracy despite a lighter network. Compared to BP-based models, Zhang et al. [42] and Kheradpisheh et al. [40] report slightly lower accuracies (88.1% and 88.0%) using IF-type neuron models with a 1000-neuron hidden layer architecture. The other competing models, Taylor et al. [39] achieve closer performance to our model employing LIFβ with 89.05% and Neftci et al. [4], performs slightly better with 89.93% accuracy and our model is very close to it.
In the CIFAR-10, our model, with a significantly larger 3072-2000-2000 architecture, achieves 51.05% accuracy using the LIF neuron model. We implemented an experiment with a similar architecture of our model based on Neftci et al. [4], utilizing the LIF neuron model, and we got a lower accuracy of 47.82%. This result suggests that while our model shows progress, the CIFAR-10 dataset remains difficult for fully connected SNNs due to training on complex, high-dimension images.
In the N-MNIST dataset, our model achieves competitive results with 96.13% accuracy using the LIF neuron model in the two hidden layers of 500 neurons architecture. Perez-Nieves et al. [41] use a single hidden layer of 200 neurons, achieving 92.70% and Taylor et al. [39] achieves 95.91% accuracy, showcasing the adaptability of our model. Neftci et al. [4], using LIFβ neurons and a similar architecture, achieves the highest accuracy at 97.47%. These results highlight strong performance for our model, although it slightly underperforms relative to Neftci et al.’s model.
The table highlights that our model demonstrates strong performance across multiple datasets, particularly excelling on MNIST with state-of-the-art accuracy while also performing very competitively on FMNIST and N-MNIST. However, for more complex datasets like CIFAR-10, where the accuracy remains relatively low at 51.05%, architectural enhancements or improved learning mechanisms are needed to address challenges posed by high-dimensional inputs. On temporal datasets like SHD, our model achieves remarkable results using a recurrent LIF architecture and outperforms other BP-based models.
Figure 3: Loss trend during training of our proposed FF-based spiking neural network model through epochs. (a) Loss changes for non-spiking datasets such as MNIST, F-MNIST, and CIFAR-10. (b) Loss changes for spiking datasets such as N-MNIST and SHD.
@@ 4.2 Experimental Results and Analysis
After comparing the performance of our model with other models based on spiking neural networks, we also conducted a comparison with some models based on the Forward-Forward algorithm that were based on artificial neural networks. Table 3 compares the performance of various FF-based models across three datasets—MNIST, F-MNIST, and CIFAR-10—using different architectures and network types. The datasets are organized into sections, and for each, the accuracy of our SNN model is compared against other referenced works that use ANN models.
In the MNIST dataset, our model, an SNN with two hidden layers containing 2000 neurons each, achieves an impressive 98.69% accuracy, outperforming or matching most ANN implementations. Ororbia et al. [46] use an ANN with the same architecture and achieve 98.66%, closely followed by Hinton’s ANN [13] with 98.63% accuracy. Gandhi et al. [45], Aminifar et al. [49], and other models with similar architectures yield accuracies ranging from 98.63% to 98.58%. Tang [43] employs a more compact ANN with two hidden layers of 200 neurons followed by a layer of 50 neurons, obtaining 98.00% accuracy. Tosato et al. [47] use a different ANN with three hidden layers of 784 neurons each, reaching a significantly lower 94.00%, demonstrating wider networks’ importance for this task.
@@
In the Fashion-MNIST dataset, our SNN-based model with two hidden layers of 2000 neurons achieves the highest accuracy of 90.27%. Ghader et al. [50] use an ANN with two hidden layers of 500 neurons, reaching a comparable 90.22%. Ororbia et al. achieve 89.60% accuracy using an ANN with the same architecture as ”our model.” Brenig et al. score 87.31%, while Tosato et al. [47] attain only 82.60% using three hidden layers of 784 neurons each. The results demonstrate the superior performance of SNNs in handling moderately complex image data.
In the more challenging CIFAR-10 dataset, our proposed SNN model with two hidden layers of 3072 neurons achieves 54.04% accuracy, and Ghader et al.’s model [50], which is our previous ANN FF-based model with two hidden layers of 500 neurons achieves 53.31%. In comparison, Brenig et al. [44] and Aminifar et al. [49] proposed the models with four hidden layers of 2000 neurons reach 47.60%, 53.95%, respectively. In general, Table 3 demonstrates that our SNN model outperforms other ANN models on MNIST and F-MNIST datasets despite its use of SNNs and relatively smaller architectures than other ANN-based approaches.
@@
One of the points that is considered in training deep neural network models is analyzing the training process. Fig. 3 shows the loss changes during the training of our FF-based spiking neural network model on both non-spiking and spiking datasets. The x-axis represents the epochs, while the y-axis represents the loss value. For the non-spiking datasets, The loss for MNIST starts at a higher value and gradually decreases, stabilizing after around 100 epochs. The Fashion-MNIST dataset follows a similar pattern but reaches a slightly lower final loss than MNIST. The CIFAR-10 dataset shows more fluctuations, indicating that the model struggles more with this dataset compared to the others. For the spiking datasets, The loss for N-MNIST starts at a moderately low value and decreases gradually, stabilizing around -0.3. The SHD dataset exhibits a much steeper initial loss decrease and stabilizes at a lower loss value compared to N-MNIST. By evaluating the overall changes in the loss values obtained in model training, it appears that the presented model has a good learning ability and has been able to converge after performing certain epochs. There is this point that when faced with more complex datasets such as CIFAR-10, the process of large changes is not always smooth and decreasing, which is expected to be improved by using different architectures.
@@
Considering the architectures and results obtained in Table 3, a relative comparison can be made between the number of parameters used in the models and their performance. Fig. 4 compares our SNN model and other ANN models, both based on the Forward-Forward algorithm, across the MNIST, F-MNIST, and CIFAR-10 datasets. In the MNIST dataset, our proposed SNN model, with 5.57 million parameters, ranks third in terms of the number of network parameters and has the best performance compared to other models. The ANN model that we presented earlier in the study [50] had 0.64 million parameters, which could be among the selected models if the goal is to select the best model with a smaller parameter volume. A similar situation occurs in the F-MNIST dataset, where our SNN model achieves the highest accuracy while ranking third in parameter count. Therefore, if reducing the number of parameters is a priority and a slight trade-off in accuracy is acceptable, the model used in [50] could be a viable alternative. In the CIFAR-10 dataset, our SNN model outperforms other ANN models in accuracy, and relative to the number of parameters used in the models, the ANN model that we presented earlier in the study [50] has meaningfully good conditions, but the size of the parameters of our SNN model and other models does not differ significantly. These comparisons demonstrate that our proposed model achieves the best performance among the evaluated models while maintaining a highly competitive parameter count. Notably, this superiority is attained without incurring excessive computational costs during the network training process.
@@
Another point commonly considered in spiking neural networks is the number of time steps required to train the network, as it directly affects both the computational cost and the inference latency. Longer time steps often provide more opportunities for the network to accumulate information and stabilize spike-based learning, potentially leading to higher accuracy, but this comes at the expense of increased energy consumption and longer training durations. Conversely, reducing the number of time steps can improve efficiency and speed. Still, it may compromise the quality of learning if the network does not have sufficient temporal resolution to capture important features. Inference within SNNs involves multiple feedforward computations, with each pass corresponding to a time step. These computations rely on sparse spikes rather than continuous activations. However, when using backpropagation to train SNNs, there is a tendency for spike signals to weaken over time, similar to the vanishing gradient problem in ANNs. To avoid losing important information, SNNs need a sufficient number of time steps, which in many cases, the number of time steps required is in the triple digits [24]. Nevertheless, our model utilizes 10-time steps in the training process on all mentioned datasets, which is a relatively small number compared to other SNN models, which shows that our proposed algorithm can have more optimal performance than other previously presented algorithms.

### v2

@@ 3 Proposed Method
MNIST F-MNIST K-MNIST CIFAR-10 N-MNIST SHD
Dataset (train/test) $60\text{k}/10\text{k}$ $60\text{k}/10\text{k}$ $60\text{k}/10\text{k}$ $50\text{k}/10\text{k}$ $60\text{k}/10\text{k}$ $8156/2264$
Input neurons $784$ $784$ $784$ 3072 $2312$ $700$
Dataset classes $10$ $10$ $10$ $10$ $10$ $20$
Epochs $300$ $300$ $300$ $300$ $300$ $500$
Learning rate $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
Batch size $4096$ $4096$ $4096$ $4096$ $4096$ $4096$
Time steps $T$ $10$ $10$ $10$ $10$ $10$ $10$
$\alpha$ (in loss) $0.6$ $0.6$ $0.6$ $0.6$ $0.6$ $0.6$
Neurons’ threshold $1.0$ $1.0$ $1.2$ $1.2$ $1.0$ $5.0$
Neurons’ decay rate $0.99$ $0.99$ $0.99$ $0.8$ $0.9$ $0.9$
4 Experiments
To evaluate the effectiveness of the proposed approach, we conducted a series of experiments. This section is organized into two parts. First, the Experimental Setup outlines our study’s datasets and implementation details. Then, the Experimental Results and Analysis section reports the performance outcomes, compares with baseline approaches, and discusses key findings and insights derived from the results.
4.1 Experimental Setup
The experiments were conducted using a PyTorch-based implementation [31], leveraging the snntorch library [32] for spiking neuron models and surrogate gradients. To evaluate the proposed method, six datasets were utilized: MNIST [33], Fashion-MNIST (F-MNIST) [34], Kuzushiji-MNIST (K-MNIST) [35], CIFAR-10 [36], Neuromorphic-MNIST (N-MNIST) [37], and Spiking Heidelberg Digits (SHD) [38]. The MNIST dataset comprises 70,000 grayscale images of hand-written digits, each 28×28 pixels wide. It is divided into 60,000 samples for training and 10,000 for testing, covering digit classes from 0 to 9. F-MNIST and K-MNIST are both drop-in replacements for the original MNIST dataset. They offer more challenging classification tasks while maintaining the same 28×28 grayscale image format and 10-class structure. F-MNIST consists of images of clothing items such as shirts, shoes, and bags, aiming to test visual recognition models beyond simple digits. K-MNIST, on the other hand, features cursive Japanese kana characters derived from historical texts, introducing additional complexity due to the intricate and variable nature of the script. The CIFAR-10 dataset includes 60,000 color images, each sized 32×32 pixels with three color channels. It spans 10 object classes, with 50,000 samples allocated for training and 10,000 for testing. Unlike MNIST and Fashion-MNIST, which use grayscale images, CIFAR-10 focuses on more visually intricate static 2D color images. The N-MNIST dataset is derived from the MNIST dataset by applying a spiking conversion. This process involves recording digit samples displayed on an LCD monitor using an ATIS sensor mounted on a motorized pan-tilt unit, which introduces motion-based neuromorphic features while retaining the same classes and sample count as MNIST. Lastly, the SHD dataset is used, which includes high-quality studio recordings of spoken digits in both German and English. The original dataset contains 10,420 samples, which are divided into 8,156 training and 2,264 testing samples spanning 20 classes.
4.2 Experimental Results and Analysis
Our experimental results and comparisons are presented in two separate tables. Table 2 provides a comparison between our FF-based SNN model and other existing SNN models, while Table 3 focuses on comparing our SNN model with various FF-based ANN implementations. Table 2 comprehensively compares spiking neural network models across five datasets: MNIST, F-MNIST, K-MNIST, N-MNIST, and SHD. The key columns detail the model name, architecture, neuron model type, and corresponding accuracy percentage, highlighting both our proposed model and various state-of-the-art models with a fully connected architecture.
In the MNIST dataset, our model achieves an impressive 98.34% accuracy using a LIFβ[39] neuron model and a two-hidden-layer architecture with 500 neurons, outperforming other models. Compared to the FF-based models, Ororbia et al. [27] use a deeper model with two hidden layers of 7000 neurons, achieving 97.46% with LIF neurons, and Terres et al. [28] achieve lower accuracy with 93.27%. Competing BP-based models from Neftci et al. [4], Comsa et al. [40], and Taylor et al. [41] use slightly different architectures and neuron models like IF or LIFβ, achieving comparable accuracies 97.87%, 97.90%, and 97.91%, respectively. However, our model outperforms these models, showcasing an improvement, potentially due to the specific architecture and neuron model used.
@@ 4.2 Experimental Results and Analysis
Table 2: Performance of our SNN model compared to the existing SNN models (* denotes self-implementation, denotes data augmentation, and β denotes learnable membrane time constants).
@@
F-MNIST Proposed model 784-500-500 LIFβ FF 89.36
@@
Perez-Nieves et al. [43] 784-200-10 LIF BP 82.20
K-MNIST Proposed model 784-500-500 LIFβ FF 91.26
Ororbia. [27] 784-7000-7000 LIF FF 91.14
Terres et al. [28] 784-1400-1400 LIF FF 85.91
@@
In the F-MNIST, our model achieves an accuracy of 89.36% with the same two hidden layers of 500 neurons architecture and LIFβ neuron model. Compared to an FF-based model, Terres et al. [28] use two hidden layers of 1400 neurons, achieving 85.68%, which our model achieves higher accuracy despite a lighter network. Compared to BP-based models, Zhang et al. [44] and Kheradpisheh et al. [42] report slightly lower accuracies (88.1% and 88.0%) using IF-type neuron models with a 1000-neuron hidden layer architecture. The other competing models, Taylor et al. [41] achieve closer performance to our model employing LIFβ with 89.05% and Neftci et al. [4], perform slightly better with 89.93% accuracy, and our model is very close to it.
The K-MNIST results further reinforce the strength of the proposed model, which achieves an accuracy of 91.26%, exceeding Ororbia’s [27] significantly denser FF-based SNN model that employs 7000 neurons in each of its two hidden layers of architecture and achieves 91.14%. This comparison highlights the efficiency of the proposed architecture, which utilizes only two hidden layers with 500 neurons each. Moreover, the proposed model substantially outperforms the model by Terres et al. [28] with the accuracy of 85.91%, which also uses the FF approach but with a wider architecture of two hidden layers with 1400 conventional LIF neurons each.
In the N-MNIST dataset, our model achieves competitive results with 97.26% accuracy using the LIF neuron model in the two hidden layers of 500 neurons architecture. Perez-Nieves et al. [43] use a single hidden layer of 200 neurons, achieving 92.70%, and Taylor et al. [41] achieve 95.91% accuracy, showcasing the adaptability of our model. Neftci et al. [4], using LIFβ neurons and a similar architecture, achieves the highest accuracy at 97.47%. These results highlight strong performance for our model, although it slightly underperforms relative to Neftci et al.’s model.
The table highlights that our model demonstrates strong performance across multiple datasets, particularly excelling on MNIST with state-of-the-art accuracy while also performing very competitively on FMNIST and N-MNIST, and on temporal datasets like SHD, our model achieves remarkable results using a recurrent LIF architecture and outperforms other BP-based models.
Table 3: Performance of our SNN model compared to the existing ANN implementation of the Forward-Forward algorithm
@@
CIFAR-10 Proposed model 3072-2000-2000 SNN 54.03
@@
After comparing the performance of our model with other models based on spiking neural networks, we also conducted a comparison with some models based on the Forward-Forward algorithm that were based on artificial neural networks. Table 3 compares the performance of various FF-based models across three datasets—MNIST, F-MNIST, and CIFAR-10—using different architectures and network types. The datasets are organized into sections, and for each, the accuracy of our 2-layer SNN model is compared against other referenced works that use ANN models. The number of neurons in the layers of our architectures is selected based on commonly reported architectures for each dataset.
Figure 4: Loss trend during training of our proposed FF-based spiking neural network model through epochs. (a) Loss changes for non-spiking datasets such as MNIST, F-MNIST, and K-MNIST. (b) Loss changes for spiking datasets such as N-MNIST and SHD.
@@
On the MNIST dataset, the proposed SNN model achieves a peak accuracy of 98.69%, surpassing compared ANN-based models, including many with substantially deeper architectures. Remarkably, it marginally outperforms the models of Ororbia et al. [48], Hinton [13], Gandhi et al. [47], and Aminifar et al. [51], each of which reports an accuracy of 98.66%, 98.63%, 98.63%, and 98.60%, respectively. This is achieved despite the proposed model employing a comparatively shallower architecture, consisting of two hidden layers with 2000 neurons each, as opposed to the four hidden layers used in the models by Hinton [13], Gandhi et al. [47], and Aminifar et al. [51].
For the F-MNIST dataset, which presents a more challenging classification task than MNIST due to greater intra-class variability, the proposed SNN model once again achieves the highest accuracy among existing ANN-based models trained using the FF algorithm, reaching 90.27%. This performance slightly surpasses that of the model proposed by Ghader et al. [52], which achieves 90.22%, and significantly outperforms other FF-trained ANN models, including those of Ororbia et al. [48], Brenig et al. [46], and Tosato et al. [49], which report accuracies of 89.60%, 87.31%, and 82.60%, respectively. These findings underscore the robustness of the proposed model across varying data distributions and its capacity to generalize effectively in classification scenarios.
The CIFAR-10 dataset serves as a more complex benchmark, consisting of natural, non-spiking image data with high-dimensional color inputs. While the increased task difficulty leads to a reduction in overall accuracy, the proposed SNN model achieves a performance of 54.03%, the highest among the compared ANN-based FF models. It slightly outperforms the deep ANN model introduced by Aminifar et al. [51], which achieves 53.95%, and significantly exceeds the results reported by Ghader et al. [52] and Brenig et al. [46], which obtain 53.31% and 47.60%, respectively. Considering the non-spiking nature and high-dimensional input of CIFAR-10, the ability of the proposed SNN to match or surpass ANN-based counterparts represents a notable achievement and further demonstrates the potential of the forward-forward learning paradigm in complex visual recognition tasks. These findings suggest that the proposed SNN architecture, when trained with the Forward-Forward learning rule, not only matches but often exceeds the performance of conventional ANN-based Forward-Forward models across diverse datasets. Its efficiency, driven by a relatively shallow design and biologically inspired neuron dynamics, points to the promise of SNNs as viable and adaptable alternatives to standard deep learning frameworks, particularly in contexts where backpropagation-free learning is desirable.
@@
One of the points that is considered in training deep neural network models is analyzing the training process. Figure 4 shows the loss changes during our FF-based spiking neural network model training on both non-spiking and spiking datasets. The x-axis represents the epochs, while the y-axis represents the loss value. For the non-spiking datasets, the loss for MNIST starts at a value and gradually decreases, stabilizing after around 100 epochs. The Fashion-MNIST dataset follows a similar pattern but reaches a slightly lower final loss than MNIST. The CIFAR-10 dataset shows more fluctuations, indicating that the model struggles more with this dataset compared to the others. For the spiking datasets, the loss for N-MNIST starts at a moderately low value and decreases gradually, stabilizing around -0.3. The SHD dataset exhibits a much steeper initial loss decrease and stabilizes at a lower loss value compared to N-MNIST. By evaluating the overall changes in the loss values obtained in model training, it appears that the presented model has a good learning ability and has converged after performing certain epochs. There is this point that when faced with more complex datasets such as CIFAR-10, the process of large changes is not always smooth and decreasing, which is expected to be improved by using different architectures.
@@
Considering the architectures and results obtained in Table 3, a relative comparison can be made between the number of parameters used in the models and their performance. Figure 5 compares our SNN model and other ANN models, both based on the Forward-Forward algorithm, across the MNIST, F-MNIST, and CIFAR-10 datasets. In the MNIST dataset, our proposed SNN model, with 5.57 million parameters, ranks third in terms of the number of network parameters and has the best performance compared to other models. The ANN model that we presented earlier in the study [52] had 0.64 million parameters, which could be among the selected models if the goal is to select the best model with a smaller parameter volume. A similar situation occurs in the F-MNIST dataset, where our SNN model achieves the highest accuracy while ranking third in parameter count. Therefore, if reducing the number of parameters is a priority and a slight trade-off in accuracy is acceptable, the model used in [52] could be a viable alternative. On the CIFAR-10 dataset, our SNN model achieves higher accuracy than all compared ANN models. In terms of parameter efficiency, the ANN model we introduced earlier in this study [52] offers a favorable balance between performance and model size. However, our proposed SNN model, with approximately 10.14 million parameters, uses significantly fewer parameters than the other evaluated models, which average around 18.14 million. These comparisons highlight that our model not only delivers the best performance among the evaluated approaches but also maintains a competitive parameter count.
@@
Another important consideration in spiking neural networks is the number of time steps required to train the network, as this parameter directly affects both computational cost and inference latency. Longer time steps typically allow the network to accumulate more information and stabilize spike-based learning, potentially leading to higher accuracy; however, this comes at the expense of increased energy consumption and prolonged training time. Conversely, reducing the number of time steps can enhance efficiency and inference speed, though it may compromise learning performance if the network lacks sufficient temporal resolution to capture salient features. Inference in SNNs involves multiple feedforward computations, each corresponding to a time step, and relies on discrete and sparse spike activity rather than continuous activations. When backpropagation is used for training, spike signals often weaken over time, resembling the vanishing gradient problem observed in traditional ANNs. To prevent the loss of important information, SNNs frequently require a relatively large number of time steps, often in the triple digits [24]. In contrast, our model employs only 10 time steps during training across all evaluated datasets. This relatively small number of time steps, compared to the majority of other SNN models, highlights the efficiency of our proposed algorithm and its ability to achieve competitive performance while maintaining a significantly lower computational burden in terms of the number of time steps.